# Med3DVLM: An Efficient Vision-Language Model for 3D Medical Image Analysis

### v1

@@ 4.4.1 Impact of Vision Encoder and Contrastive Loss
We evaluated the effect of different vision encoders (ViT3D vs. DCFormer) and contrastive loss functions (CLIP vs. SigLIP) on image-text retrieval performance. The results in Table 5 indicate that DCFormer significantly outperforms ViT3D, particularly at larger test set sizes, confirming its effectiveness in preserving spatial and structural information in 3D medical images. Additionally, switching from CLIP to SigLIP further improves retrieval performance, where Med3DVLM achieves the highest recall across all test sizes. These findings suggest that SigLIP better handles small-batch medical datasets by improving semantic alignment and feature stability in contrastive learning.
Table 5: The impact of vision encodes and loss functions on image-text retrieval. DCFormer-S: small version of DCFormer.
Methods ViT3D CLIP DCFormer CLIP DCFormer-S SigLIP (Ours)
Test Samples 100 500 1000 2000 100 500 1000 2000 100 500 1000 2000
IR R@1 63.00 33.40 25.90 19.10 84.00 52.00 41.10 31.20 92.00 77.20 69.30 61.00
R@5 89.00 67.80 57.20 44.70 95.00 85.40 77.20 67.15 95.00 92.80 91.40 87.20
R@10 95.00 81.00 71.60 59.35 97.00 90.80 85.90 78.35 96.00 95.20 93.90 91.25
TR R@1 57.00 32.00 25.50 19.40 81.00 53.20 41.10 30.35 90.00 78.40 71.30 63.90
R@5 90.00 67.40 56.90 45.35 95.00 84.80 76.40 64.70 95.00 93.40 91.40 88.05
R@10 94.00 81.80 70.70 59.10 98.00 91.80 86.00 78.20 95.00 95.60 94.10 91.90
@@
Method BLEU ROUGE METEOR BERTScore
2xMLP 15.63 19.86 15.10 84.24
2xMLP-H 15.99 20.25 15.68 84.31
1xMLP-Mixer-H 23.93 27.70 23.25 85.73
2xMLP-Mixer-H 36.88 40.25 36.42 88.11
4.5 Impact of Multi-Modal Projectors
We compared the performance of our model configured with different multi-modal projectors. The results in Table 6 show that DCFormer with 2×MLP-Mixer achieves the highest BLEU, ROUGE, and METEOR scores, confirming that this architecture enables better representation fusion for text generation. In contrast, replacing the 2×MLP-Mixer with simpler MLP-based projection layers leads to a significant drop in fluency and content relevance, as evidenced by lower METEOR scores. These results highlight the importance of low-high feature integration for structured radiology report generation.
Additionally, we investigated the role of multi-modal projectors in both open-ended and closed-ended VQA. In open-ended VQA, where generating clinically relevant responses requires deeper contextual understanding, the 2×MLP-Mixer projector achieves the highest BLEU, ROUGE, and METEOR scores (Table 7). Meanwhile, models using solely MLP-based projectors generate reasonable responses but show limitations in coherence and fluency, reinforcing the importance of multi-level feature aggregation in medical VQA. For closed-ended VQA, the performance gap between projectors is smaller, suggesting that classification-based answering relies more on direct pattern recognition and dataset priors rather than feature expressiveness. Nonetheless, the 2×MLP-Mixer maintains the highest overall accuracy, indicating that richer feature fusion still contributes to more precise and consistent decision-making.
@@ 5.3 Challenges and Future Directions
As future directions, incorporating structured clinical knowledge could help mitigate hallucinations and improve interpretability. Validating Med3DVLM across diverse clinical datasets and imaging protocols will be essential to ensure its reliability in real-world settings. Moreover, extending the model to support uncertainty-aware response generation may further enhance its trustworthiness in clinical decision support.

### v2

@@ 4.4.2 Impact of Vision Encoder and Contrastive Loss
Table 8: The impact of different vision encoders and loss functions on image-text retrieval. DCFormer-S represents small version of DCFormer.
Methods ViT3D CLIP ViT3D SigLIP
Test Samples 100 500 1000 2000 100 500 1000 2000
IR R@1 63.00 33.40 25.90 19.10 75.00 55.40 46.20 38.65
R@5 89.00 67.80 57.20 44.70 91.00 84.40 79.50 70.50
R@10 95.00 81.00 71.60 59.35 94.00 89.80 87.10 79.45
TR R@1 57.00 32.00 25.50 19.40 78.00 59.60 49.40 39.70
R@5 90.00 67.40 56.90 45.35 91.00 84.20 77.40 69.20
R@10 94.00 81.80 70.70 59.10 93.00 88.80 85.50 78.40
Methods DCFormer-S CLIP DCFormer-S SigLIP
Test Samples 100 500 1000 2000 100 500 1000 2000
IR R@1 76.00 51.80 40.80 29.85 92.00 77.20 69.30 61.00
R@5 94.00 86.00 74.60 64.60 95.00 92.80 91.40 87.20
R@10 97.00 91.40 84.90 75.95 96.00 95.20 93.90 91.25
TR R@1 83.00 52.60 43.50 32.60 90.00 78.40 71.30 63.90
R@5 94.00 86.00 75.00 64.10 95.00 93.40 91.40 88.05
R@10 98.00 91.60 85.70 76.70 95.00 95.60 94.10 91.90
We evaluated the impact of different vision encoders (ViT3D vs. DCFormer-S) and contrastive loss functions (CLIP vs. SigLIP) on image-text retrieval. As shown in Table 8, both image retrieval (IR) and text retrieval (TR) benefit significantly from using DCFormer-S over ViT3D, particularly at larger test sizes. For example, with 2,000 test samples and SigLIP loss, DCFormer-S achieves an R@1 of 61.00% for IR and 63.90% for TR, compared to 38.65% and 39.70% respectively with ViT3D.
Across all encoder types and test sizes, switching from CLIP to SigLIP consistently improves performance. For example, with the ViT3D encoder and 2,000 test samples, SigLIP improves IR R@1 from 19.10% to 38.65%, and TR R@1 from 19.40% to 39.70%. With the DCFormer-S encoder, the gains are also evident: IR R@1 increases from 29.85% to 61.00%, and TR R@1 from 32.60% to 63.90%. These improvements are especially pronounced at larger evaluation scales, where SigLIP enhances alignment stability without requiring large negative batches.
@@
Method BLEU ROUGE METEOR BERTScore Parameters Flops
2xMLP 15.63 19.86 15.10 84.24 15.60 M 0.50 G
2xMLP-H 15.99 20.25 15.68 84.31 16.98 M 4.14 G
1xMLP-Mixer-H 23.93 27.70 23.25 85.73 29.91 M 3.85 G
2xMLP-Mixer-H 36.88 40.25 36.42 88.11 47.22 M 6.14 G
4.4.3 Impact of Multi-Modal Projectors
@@
We evaluated the impact of different multi-modal projector designs on radiology report generation, open-ended VQA, and closed-ended VQA. The 2×MLP variant consists of a simple two-layer multilayer perceptron applied to global visual features, using a linear activation followed by GELU and another linear transformation. The 2×MLP-H variant incorporates hierarchical features from the final and penultimate layers of the vision encoder; each feature map is independently processed by a 2×MLP block, and their outputs are concatenated. The 1×MLP-Mixer-H design replaces the MLPs with a single MLP-Mixer block that jointly processes the concatenated hierarchical features, enabling token and channel mixing. The 2×MLP-Mixer-H configuration, used in our final model, stacks two MLP-Mixer blocks and includes an additional hidden projection layer to further enhance multi-scale feature interaction. To provide a complete view of computational efficiency, we also report the number of parameters and FLOPs for each projector variant in Table 9
[…]
@@ 5.3 Challenges and Future Directions
While the M3D dataset provides a large-scale and diverse benchmark, it primarily consists of CT images and focuses on a limited set of tasks including retrieval, report generation, and VQA. To ensure broader generalization and real-world applicability, future evaluation should incorporate other imaging modalities such as MRI and ultrasound, as well as datasets that differ in patient demographics, clinical settings, and scanner protocols. In addition, extending the scope of evaluation to include tasks such as medical image synthesis, modality translation, and segmentation would offer a more comprehensive assessment of the model’s robustness and clinical utility across a wider range of scenarios.
As for future directions, incorporating structured clinical knowledge could help reduce hallucinations and improve interpretability. Furthermore, as noted above, validating Med3DVLM across diverse clinical datasets and imaging protocols will be critical to ensuring its reliability in real-world settings. Expanding the model to support uncertainty-aware response generation may also enhance its trustworthiness in clinical decision support.
5.4 Computational Complexity
We compare Med3DVLM and M3D-LaMed in terms of parameter count and FLOPs, as summarized in Table 12. Med3DVLM replaces the 3D Vision Transformer in M3D-LaMed with the DCFormer encoder, reducing the vision backbone from 87.4M parameters and 253.23G FLOPs to 18.2M parameters and 21.59G FLOPs. Although Med3DVLM uses a larger multi-modal projector (47.2M vs. 19.9M parameters), this overhead is minor compared to the efficiency gains from the encoder. The total model size increases slightly (7.6B vs. 6.9B) due to a larger LLM, which is fine-tuned efficiently using LoRA.
We do not report direct comparisons of inference time and memory usage, as M3D-LaMed includes an additional segmentation module not present in our model. This makes runtime comparisons non-equivalent. To ensure fairness, we focus on theoretical complexity metrics that isolate the vision-language components.
Table 12: Comparison of parameters and FLOPs for each module in Med3DVLM. M3D-LaMed values are reported from the original paper.
Module M3D-LaMed Med3DVLM
Params FLOPs Params FLOPs
3D Image Encoder 87.4M 253.23G 18.2M 21.59G
Multimodal Projector 19.9M 5.10G 47.2M 6.14G
LLM with LoRA 6.7B - 7.6B -
All 6.9B - 7.6B -